OpenStudy (anonymous):

guys is my answer right? thankyou (x^3-27)/(x^3+64) times (x+4)/(x2+3x+9) =(x-3)(x^2+3x+9)/(x+4)(x^2-4x+16) times (x+4)/(x^2+3x+9 =(x+4)/x^23x+9 is this right?? thankyou guys ^^ takecare

OpenStudy (perl):

take care? where you going

hartnn (hartnn):

what did u do after this step ? (x-3)(x^2+3x+9)/(x+4)(x^2-4x+16) times (x+4)/(x^2+3x+9 x+4 gets cancelled, isn't it ?

hartnn (hartnn):

and (x^2+3x+9)

OpenStudy (anonymous):

\[x-3 \]

hartnn (hartnn):

so what remains is (x-3) / (x^2-4x+16)

OpenStudy (anonymous):

@hartnn yes i cancelled it

OpenStudy (anonymous):

@perl LOL.

OpenStudy (perl):

:)

OpenStudy (anonymous):

@hartnn yes. is it wrong? @curiousshubham . is the ans x-3?

OpenStudy (anonymous):

@xtel yep!

hartnn (hartnn):

x+4 gets cancelled , your answer --->(x+4)/x^2+3x+9 are the terms that actually gets cancelled so what remains is (x-3) / (x^2-4x+16)

hartnn (hartnn):

@curiousshubham where did (x^2-4x+16) go ?

OpenStudy (anonymous):

@hartnn is cancelling the nos. is the proper way ? thankyou

OpenStudy (anonymous):

@xtel \[\frac{ x^{3}-27 }{ x^{3}+64 }\times \frac{ x+4 }{ x^{2}+3x+9 }\] This is your question?

OpenStudy (anonymous):

@curiousshubham yes, thats it

OpenStudy (anonymous):

Sorry @hartnn I left \[x^{2}-4x+16\] @xtel the solution is \[\frac{ x-3 }{ x^{2}-4x+16 }\]

hartnn (hartnn):

no problem.

OpenStudy (anonymous):

@curiousshubham and @hartnn thankyou guys :))

hartnn (hartnn):

welcome :)

OpenStudy (anonymous):

you are welcomed...

OpenStudy (anonymous):

:)